# Framing in the Presence of Supporting Data: A Case Study in U.S. Economic News

### v2

@@ A.8 Additional Framing Analysis Graphs
Figure 13: Selection of economic indicators referencing jobs numbers from 2015 through 2023 in the New York Times, Washington Post, and Wall Street Journal
@@
Figure 14: Framing of articles referencing job numbers from 2015 through 2023 in the New York Times
Figure 15: Framing of articles referencing job numbers from 2015 through 2023 in the Washington Post
Figure 16: Framing of articles referencing job numbers from 2015 through 2023 in the Wall Street Journal
Figures 14, 15, and 16, show how the New York Times, the Washington Post and the Wall Street Journal use jobs indicators in their reporting.
@@
Figure 17: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the New York Times
Figure 18: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the Washington Post
Figure 19: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the Wall Street Journal
Figures 17, 18, and 19, show how the New York Times, the Washington Post, and the Wall Street Journal use price (including energy price) indicators in their reporting. Once again, the New York Times and Washington Post display consistently negative framing of price indicators, while the Wall Street Journal has a more balanced framing of such indicators. It is interesting that the framing of price indicators in the New York Times and Washington Post seems almost static, with no obvious change following the rise in inflation. This differs from the shifts in framing we see in both papers in Figures 14 and 15.

### v3

@@ A.10 Additional Framing Analysis Graphs
Figure 15: Selection of economic indicators referencing jobs numbers from 2015 through 2023 in the New York Times, Washington Post, and Wall Street Journal. Aggregated Quarterly. Monthly payroll data can be seen in the dotted blue line.
@@
Figure 16: Framing of articles referencing job numbers from 2015 through 2023 in the New York Times. Aggregated Quarterly. Monthly payroll data can be seen in the dotted black line.
Figure 17: Framing of articles referencing job numbers from 2015 through 2023 in the Washington Post. Aggregated Quarterly. Monthly payroll data can be seen in the dotted black line.
Figure 18: Framing of articles referencing job numbers from 2015 through 2023 in the Wall Street Journal. Aggregated Quarterly. Monthly payroll data can be seen in the dotted black line.
Figures 16, 17, and 18, show how the New York Times, the Washington Post and the Wall Street Journal use jobs indicators in their reporting.
@@
Figure 19: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the New York Times. Aggregated Quarterly. Monthly CPI data can be seen in the dotted black line.
Figure 20: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the Washington Post Aggregated Quarterly. Monthly CPI data can be seen in the dotted black line.
Figure 21: Framing of articles referencing price (price & energy) numbers from 2015 through 2023 in the Wall Street Journal. Aggregated Quarterly. Monthly CPI data can be seen in the dotted black line.
Figures 19, 20, and 21, show how the New York Times, the Washington Post, and the Wall Street Journal use price (including energy price) indicators in their reporting. Once again, the New York Times and Washington Post display consistently negative framing of price indicators, while the Wall Street Journal has a more balanced framing of such indicators. It is interesting that the framing of price indicators in the New York Times and Washington Post seems almost static, with no obvious change following the rise in inflation. This differs from the shifts in framing we see in both papers in Figures 16 and 17.